4. (35) Suppose we have the following recursive function in C++: a. (5) What is the returned value of fib(5)? b. (10) Supp
Answer #1

Answer a)The value returned by the function fib(5) is 5 From the definition fib(n-1)+fib(n-2) fib(5)=fib(4-1)+fib(4-2)=5 b) F